ml 0xygen consumed/ 1 min x 60 minutes/ 1 hr = ml 0xygen /hr
my question is 7 ml oxygen consumed / 1 min x 60 minutes / 1 hr = ml 0xygen/ hr

Answer by stanbon(75887)

[ml 0xygen consumed/ 1 min] x (60 minutes/ 1 hr) = [(60ml 0xygen) /hr]
-----------------------------

my question is 7 ml oxygen consumed / 1 min x 60 minutes / 1 hr = ml 0xygen/ hr
----
7 (ml/min) * (60min/hr) = (7*60)(ml/hr) = 420(ml/hr)
----------------
Cheers,
Stan H.
-------------